OpenStudy (anonymous):

Find the quantity indicated. r varies directly as s. Find r when s = 4 and k = 15.

jimthompson5910 (jim_thompson5910):

"r varies directly as s" so this means r = k*s where k is some constant

jimthompson5910 (jim_thompson5910):

The next bit "Find r when s = 4 and k = 15" allows us to go from r = k*s to r = 15*4 r = 60

OpenStudy (anonymous):

Thank you. You explained it as simple s possible and now I understand.

jimthompson5910 (jim_thompson5910):

you're welcome